OpenStudy (anonymous):

And AGAiN.. 3x+2y=-17 2x-5y=14

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

solve for linear equation by substituting.

OpenStudy (anonymous):

2y = -17 - 3x y = (-17 - 3x )/2 2x -5[(-17 - 3x )/2] = 14 2x - (-58 - 15x) / 2 = 14 4x + 58 +15x = 28 19x = -30 x = -30/19 ?

OpenStudy (anonymous):

y = [-17 -3(-30/19)]/2 y = (-17 - 90/19) / 2 y = ( -323/19 - 90/19 ) / 2 y = (-413/19) / 2 y = -419 / 38

OpenStudy (anonymous):

Multiply the first equation by 2 and the second by 3 to get\[6x + 4y = -34\] and\[6x -15y = 42\] Now you could just subtract one from the other to make the 6x go away but the question says substitution so from the second one\[6x = 42 + 15y\] and put this into the first to get\[42 + 15y +4y = -34\] so \[19y = -76, y=-76/19 =-4\] Now put this value of y into one of the original equations to get the value of x. Then put the value of x and y into one of the equations to check your work is right. @josephniel14 you made a small mistake in your calculation for x. 85 not 58.

OpenStudy (anonymous):

@estudier: Right. Sorry. LOL. I don't know why but I mistakely do that all the time.